OpenStudy (anonymous):

[2.08] The formula for degrees Celsius is: C =5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 8 degrees Celsius is

OpenStudy (anonymous):

Isn't the degree for Celsius: C = 5/9(F − 32)

OpenStudy (anonymous):

@stgreen you were right, unless she posted the question wrong. @ahoward79 could you check the question?

OpenStudy (anonymous):

Oh you are right, posted it wrong

OpenStudy (anonymous):

@jazy yeah...i deleted when i came to know the true one

OpenStudy (anonymous):

Can someone show me how to do it with the correct formula?

OpenStudy (anonymous):

C=5/9(F-32) 9C=5F-160 5F=9C+160 F=9/5(C+32)